OpenStudy (anonymous):

Add & simplify: x+1/x + x-3/3x :) /= over

OpenStudy (anonymous):

I will answer your question as I think you intended it, instead of exactly as you wrote it.... I think the problem is\[\frac{ x+1}{x} +\frac{ x-3}{3x}=\frac{3x+3+x-3}{3x}=\frac{4x}{3x}=\frac{4}{3}\]

OpenStudy (anonymous):

Find the common denominator, combine like terms, simplify. Straightforward algebra. Got it?